separate the differential equation into variables

(x2-1)y,+2xy2=0


1
Expert's answer
2022-02-01T17:27:32-0500

"(x^2 -1)y'+2xy^2=0\\
(x^2 -1)dy+2xy^2dx=0\\
(x^2 -1)dy=-2xy^2dx\\
\frac{dy}{y^2}=\frac{-2xdx}{x^2-1}\\
\int\frac{dy}{y^2}=\int\frac{-2xdx}{x^2-1}\\
-\frac{1}{y}=-\ln|x^2-1|-C\\
\frac{1}{y}=\ln|x^2-1|+C\\"